OpenStudy (anonymous):

Solve -3(2x - 9) = -3. -4 4 -5 5

OpenStudy (shaik0124):

-6x+27=-3 -6x+27+3=-3+3 -6x+24=0 -6x=-24 x=24/6 x=4

OpenStudy (joannablackwelder):

27+3 is 30

OpenStudy (shaik0124):

sorry -6x+30=0 -6x=-30 x=5 is answer

OpenStudy (joannablackwelder):

Yes, I get 5. Do you understand why @parisgirl1515 ?

OpenStudy (anonymous):

-6x+27=-3 -6x=-3-27 -6x=-30 x=-30/-6 x=5